# EGG-SR: Embedding Symbolic Equivalence into Symbolic Regression via Equality Graph

### v1

@@ 4 Related Works
Symbolic Equivalence is a core concept in program synthesis and mathematical reasoning (DBLP:journals/pacmpl/WillseyNWFTP21). In SQL query optimization, it enables rewriting queries into equivalent but more efficient forms (DBLP:journals/corr/abs-2407-12794). In hardware synthesis, it supports cost-aware rewrites such as optimized matrix multiplications (DBLP:conf/fccm/UstunSYYZ22). In formal methods, it accelerates automated theorem proving via normalization and equivalence checking (DBLP:journals/pacmpl/KurashigeJGBNIJ24). In mathematical language processing, it enables the generation of paraphrased forms of mathematical expressions (DBLP:conf/naacl/ZhengWGK25). In symbolic regression, DBLP:conf/gecco/FrancaK23 leverages e-graphs to mitigate over-parameterization of candidate expressions. DBLP:conf/gecco/olivetti25 incorporates e-graphs into genetic programming to detect and eliminate redundant individuals. In contrast, our approach employs e-graphs not only for simplified variants but also to generate a richer set of equivalent variants, enabling easy integration into many symbolic regression algorithms.
Knowledge-Guided Scientific Discovery. Recent efforts have explored incorporating physical and domain-specific knowledge to accelerate the symbolic discovery process. AI-Feynman (journal/2020/aifrynman; DBLP:conf/nips/UdrescuTFNWT20; keren2023computational; cornelio2023combining) constrained the search space to expressions that exhibit compositionality, additivity, and generalized symmetry. Similarly, DBLP:journals/corr/abs-2303-03192 encoded physical unit constraints into learning to eliminate physically impossible solutions. Other works further constrained the search space by integrating user-specified hypotheses and prior knowledge, offering a more guided approach to symbolic regression (DBLP:conf/icml/BendinelliBK23; kamienny:tel-04391194; DBLP:journals/iclr/llmsr25; taskin2025knowledge). DBLP:conf/iclr/ZhangCXB025 propose using retrieval-augmented generation to acquire new, informative data that accelerates the overall learning process. Our Egg presents a new idea in knowledge-guided learning that is orthogonal to existing approaches. An interesting direction for future work is to investigate whether Egg can be combined with these knowledge-guided strategies to further improve symbolic discovery.
Equivalence-aware Learning. Equivalence and symmetry have long been recognized as crucial for improving efficiency in search and learning. In MCTS, transposition tables exploit graph equivalence by merging nodes that represent the same underlying state, avoiding redundant rollouts and accelerating convergence (DBLP:conf/cig/ChildsBK08). More recent extensions explicitly leverage symmetries to prune symmetric branches of the search space (DBLP:journals/kbs/SaffidineCM12; DBLP:conf/acml/LeurentM20). In reinforcement learning, symmetries in the state–action space have been used to accelerate convergence (grimm2020value). Large Language Models (LLMs) also benefit from equivalence-awareness, particularly in code generation (sharma2025assessing).
@@ 5 Experiments
We show that (1) Egg enhances existing learning algorithms in discovering expressions with smaller Normalized MSEs. (2) In case studies, Egg consistently exhibits both time and space efficiency.
5.1 Experiment Settings
Choice of Datasets We consider the Trigonometric dataset (DBLP:conf/pkdd/JiangX23), where each group contains 10 randomly sampled expressions. Also, we select 10 challenging equations from the Feynman dataset (DBLP:conf/nips/UdrescuTFNWT20). In terms of baselines,
Evaluation Metrics. The goodness-of-fit indicates how well the learning algorithms perform in discovering underlying expressions. We use the normalized mean-squared error (NMSE) of the best-predicted expression by each algorithm, on a separately-generated testing dataset. Given a testing dataset ${D}_{\text{test}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ generated from the ground-truth expression, we measure the goodness-of-fit of a predicted expression $\phi$, by evaluating the normalized-mean-squared-error (NMSE):
$\text{NMSE}(\phi)=\frac{1}{n\sigma_{y}^{2}}\sum_{i=1}^{n}(y_{i}-\phi(\mathbf{x}_{i}))^{2}$ (5)
The empirical variance $\sigma_{y}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}\right)^{2}}$. We use the NMSE as the main criterion for comparison in the experiments and present the results on the remaining metrics in the case studies. The main reason is that the Normalized-MSE is less impacted by the output range. The exact experimental configuration details are provided in Appendix E.
@@ 5.2 Overall Benchmarks
Impact of Egg on MCTS. We conduct two analyses to evaluate the impact of integrating Egg into standard MCTS: (1) the median normalized MSE of the TopK ($K=10$) expressions identified at the end of training, and (2) the growth of the search tree, measured by the number of explored nodes over learning iterations. The dataset is selected from (DBLP:conf/pkdd/JiangX23) as the expressions contain $\sin,\cos$ operators, which contain many symbolic-equivalence variants. The detailed experimental setups, including the rewrite rules and datasets used in each comparison, are provided in Appendix E.
@@
Table 1 shows that Egg-MCTS consistently discovers expressions with lower normalized quantile scores compared to standard MCTS. Figure 3(Left) illustrates that Egg-MCTS maintains a broader and deeper search tree, indicating exploration of a larger and more diverse search space. Across various datasets, augmenting MCTS with Egg improves symbolic expression accuracy. This improvement is primarily due to the effectiveness of our rewrite rules, which cover a rich set of trigonometric identities and enable efficient exploration of symbolic variants in trigonometric expression spaces.
@@
Egg-MTCS <<1E-6 <<1E-6 0.006 0.009 $\underline{0.005}$ $\underline{0.012}$ $\underline{0.091}$ $\underline{0.121}$
MTCS 0.006 0.033 0.144 0.147 $0.015$ $0.007$ $0.138$ 0.150
@@
LLM-SR (Mixtral) <<1E-6 $0.0002$ 0.0030{0.0030} 0.0291{0.0291} $\underline{0.0026}$ $\underline{0.0037}$ $0.0162$ $0.0946$
@@
Impact of Egg on LLM. We evaluate the impact of augmenting LLMs with richer feedback prompts that incorporate equivalent expressions generated by Egg. Following the dataset and experimental setup from the original paper (DBLP:journals/iclr/llmsr25), we summarize the results in Table 2. The result of LLM-SR directly uses the reported result in (DBLP:journals/iclr/llmsr25). The results show that integrating Egg enables the LLM to discover higher-quality expressions under the same experimental conditions.
@@ 5.3 Case Analysis
Space Efficiency of Egg. We evaluate the space efficiency of the e-graph representation in comparison to a traditional array-based approach. The array-based method explicitly stores each expression variant as a unique sequence, leading to exponential memory growth. In contrast, the e-graph compactly encodes multiple equivalent expressions by sharing common sub-expressions.
We benchmark the memory consumption of storing all equivalent variants of input expressions under two settings: (1) $\phi=\log(x_{1}\times\ldots\times x_{n})$, using the logarithmic identity $\log(ab)\leadsto\log a+\log b$, and (2) $\phi=\sin(x_{1}+\ldots+x_{n})$, using the trigonometric identity $\sin(a+b)\leadsto\sin(a)\cos(b)+\sin(b)\cos(a)$. Both settings yield $2^{n-1}$ equivalent variants.
Figure 5 presents the memory consumption as a function of the input variable size $n$. The results show that the e-graph representation achieves significantly lower memory usage compared to the array-based approach. Figure 5(right) visualizes two sample e-graphs, which illustrate how shared sub-expressions are stored only once, contributing to the space efficiency of Egg. The corresponding e-graphs for $n=2,3,4$ are visualized in Appendix Figures 13 and 14.
Figure 4: Egg uses less memory than the array-based approach for two settings: (Left) $\log\left(x_{1}\times\ldots x_{n}\right)$ rewritten using $\log(ab)\leadsto\log a+\log b$. (Right) $\sin\left(x_{1}+\ldots x_{n}\right)$ rewritten using $\sin(a+b)\leadsto\sin a\cos b+\sin a\cos b$.
@@
Time Efficiency of Egg with DRL. As shown in Figure 5, we benchmark the runtime of the four main computations in Egg-DRL on the selected “sincos(3,2,2)” dataset: (1) sampling sequences of rules from the sequential decoder, (2) fitting coefficients in symbolic expressions to the training data, (3) generating equivalent expressions via Egg, and (4) computing the loss, gradients, and updating the neural network parameters. We consider two settings for the sequential decoder: a 3-layer LSTM with hidden dimension 128, and a decoder-only Transformer with 6 attention heads and hidden dimension 128. The Egg module contributes minimal computational overhead relative to more expensive steps such as coefficient fitting and network updates. This highlights the practicality of incorporating Egg into DRL-based symbolic regression frameworks.
Additional E-graph Visualizations. To further demonstrate the effectiveness of the proposed Egg module, we present several additional e-graph visualizations generated using our implemented API on selected complex expressions from the Feynman dataset (journal/2020/aifrynman), each highlighting different rewrite rules (see Appendix F.2).
@@ 6 Conclusion
In this paper, we introduced Egg-DR, a unified framework that integrates symbolic equivalence into symbolic regression through equality graphs (e-graphs) to accelerate the discovery of optimal expressions. Our theoretical analysis establishes the advantages of Egg-MCTS over standard MCTS and Egg-DRL over conventional DRL algorithms. Extensive experiments further demonstrate that Egg consistently enhances the ability of existing methods to uncover high-quality governing equations from experimental data. Currently, the collection of rewrite rules relies on manual effort. In future work, we plan to develop systematic approaches for discovering more rewrite rules and to extend the proposed method to related tasks, such as code optimization and program synthesis.
Acknowledgments
This research was supported by NSF grant CCF-1918327, NSF Career Award IIS-2339844, DOE – Fusion Energy Science grant: DE-SC0024583.
Contents
Availability of Egg, Baselines and Dataset Please find our code repository at: https://www.github.com/jiangnanhugo/egg-sr
  • Egg. The implementation of our Egg module is located in the folder “
    src/equality_graph/”.
  • Datasets. The list of datasets used in our experiments can be found in
    datasets/scibench/scibench/data/”.
  • Baselines. The implementations of several baseline algorithms, along with our adapted versions, are provided in the ”algorithms/” folder. Execution scripts for running each algorithm are also included.
We also provide a README.md file with instructions for installing the required Python packages and dependencies.
We summarize the supplementary material as follows: Section A provides implementation details of the proposed Egg. Sections B and C present a detailed theoretical analysis of the proposed method. Section E describes the experimental setup and configurations. Finally, Section F presents additional experimental results.
Appendix A Implementation of Egg Module
In what follows, we present the implementation of a unified framework that employs a context-free grammar to represent symbolic expressions, rewrite rules, and e-graphs.
A.1 Context-free Grammar Definition for Symbolic Expression
The context-free grammar for symbolic regression is defined in Section 2. Figure 6 further illustrates how an expression such as $\phi=c_{1}\log(x_{1})$ can be constructed from the start symbol by sequentially applying grammar rules. We begin with the multiplication rule $A\to A\times A$, which expands the initial symbol $A$ in $\phi=A$ to yield $\phi=A\times A$. Continuing this process, each non-terminal symbol is recursively expanded using the appropriate grammar rules until we obtain the valid expression $\phi=c_{1}\log(x_{1})$.
@@ A.1 Context-free Grammar Definition for Symbolic Expression
where the loss function $\ell$ is typically the normalized mean squared error (NMSE), which is defined in Equation 5.
@@ A.2 Implementation of Rewrite Rules
class Rule:
def __init__(self, lhs, rhs):
self.lhs = rules_to_s_expr(lhs)
self.rhs = rules_to_s_expr(rhs)
Each rule encodes one direction of applying a mathematical identity. Prior research (DBLP:conf/gecco/olivetti25) has primarily used rewrite rules to simplify expressions, typically rewriting a longer expression into a shorter one. In contrast, our work considers both directions of each identity, enabling us to systematically generate the complete set of symbolically equivalent expressions.
@@
    # Commutative laws
    Rule(lhs=[’A->(A+A)’, A->a’, A->b’], rhs=[’A->(A+A)’, A->b’, A->a’])
    Rule(lhs=[’A->A*A’, A->a’, A->b’], rhs=[’A->A*A’, A->b’, A->a’])
@@
A.3 Implementation of E-Graph Functionalities
@@
class ENode:
def __init__(self, operator: str, operands: List):
self.operator, self.operands = operator, operands
[…]
# Convert to the canonical form of this ENode
[…]
def __init__(self, id):
self.id = id
self.parent = None
def find_parent(self):
# Return the representative parent ID (using union-find algorithm)
def __repr__(self):
return e-class{}’.format(self.id)
The e-graph structure is implemented in the following class. The dictionary hashcons stores all the edges from an e-node to an e-class. The API supports the key operations of (1) adding new expressions to the e-graph by calling add_enode, and (2) merging two e-classes into one:
class EGraph:
# hashcons: stores a mapping from ENode to its corresponding EClass
hashcons: Dict[ENode, EClassID] = {}
[…]
# Add an ENode into the e-graph
[…]
# Return all e-classes and their associated e-nodes
[…]
# Merge two e-classes using the union-find structure
[…]
# Construct a new expression in the e-graph from a rule RHS
Construction The e-graph is built by repeatedly applying a set of rewrite rules to the e-graph. Each rewrite rule consists of a left-hand side (LHS) and a right-hand side (RHS) expression. The function apply_rewrite_rules determines which new subgraphs to construct and where to merge them within the existing e-graph.

### v2

@@ 4 Related Works
Knowledge-Guided Scientific Discovery. Recent efforts have explored incorporating physical and domain-specific knowledge to accelerate symbolic discovery. AI-Feynman [journal/2020/aifrynman, DBLP:conf/nips/UdrescuTFNWT20, keren2023computational, cornelio2023combining] constrained the search space to expressions that exhibit compositionality, additivity, and generalized symmetry. Similarly, tenachi2023deep encoded physical unit constraints into learning to eliminate physically impossible solutions. Other works further constrained the search space by integrating user-specified hypotheses and prior knowledge, offering a guided approach to symbolic regression [DBLP:conf/icml/BendinelliBK23, kamienny:tel-04391194, DBLP:journals/iclr/llmsr25, taskin2025knowledge, DBLP:conf/iclr/ZhangCXB025]. Our Egg-SR presents a new idea in knowledge-guided learning that is orthogonal to existing approaches.
Symbolic Equivalence is a central concept in program synthesis and mathematical reasoning [DBLP:journals/pacmpl/WillseyNWFTP21]. In SQL query optimization, it rewrites queries into time-efficient forms [DBLP:journals/corr/abs-2407-12794]. In hardware synthesis, it supports cost-aware rewrites such as optimized matrix multiplication [DBLP:conf/fccm/UstunSYYZ22]. In formal methods, it accelerates automated theorem proving through normalization and equivalence checking [DBLP:journals/pacmpl/KurashigeJGBNIJ24]. In mathematical reasoning, it is used to generate paraphrases of math expressions [DBLP:conf/naacl/ZhengWGK25].
In symbolic regression, DBLP:conf/gecco/FrancaK23 leverages e-graphs to mitigate over-parameterization in candidate expressions. DBLP:conf/gecco/olivetti25, de2025equality further incorporates e-graphs into genetic programming (GP) to detect and eliminate redundant individuals, encouraging GP to explore novel expressions. A recent follow-up work [rEGGression] provides a richer API for interacting with GP. This study advances existing work by offering a unified interface for encoding known mathematical equalities as e-graphs, enabling equivalence-aware learning across several modern symbolic regression algorithms, together with theoretical guarantees.
Equivalence-aware Learning. Equivalence and symmetry have long been recognized as crucial for improving efficiency in search and learning. In MCTS, transposition tables exploit equivalence by merging nodes that represent the same underlying state, avoiding redundant rollouts and accelerating convergence [DBLP:conf/cig/ChildsBK08]. More recent extensions explicitly leverage symmetries to prune symmetric branches of the search space [DBLP:journals/kbs/SaffidineCM12, DBLP:conf/acml/LeurentM20]. In reinforcement learning, symmetries in the state–action space have been used to accelerate convergence [grimm2020value]. LLMs also benefit from equivalence-awareness, particularly in code generation [sharma2025assessing].
@@ 5 Experiments
We show that (1) Egg enhances existing learning algorithms in discovering expressions with smaller Normalized MSEs. (2) In case studies, Egg consistently exhibits both time and space efficiency. The detailed experimental setups and datasets used in each comparison, are provided in Appendix C.
@@ 5.1 Overall Benchmarks
Impact of Egg on MCTS. We conduct two analyses to evaluate the impact of integrating Egg into standard MCTS: (1) the median normalized MSE of the TopK ($K=10$) expressions identified at the end of training, and (2) The growth of the search tree, measured by the number of explored nodes over learning iterations.
@@
Table 1 shows that Egg-MCTS consistently discovers expressions with lower normalized quantile scores compared to standard MCTS. The dataset is selected from [DBLP:conf/pkdd/JiangX23] as the expressions contain $\sin,\cos$ operators, which contain many symbolic-equivalence variants.
Figure 3(Left) illustrates that Egg-MCTS maintains a broader and deeper search tree, indicating exploration of a larger and more diverse search space. Across various datasets, augmenting MCTS with Egg improves symbolic expression accuracy. This improvement is primarily due to the effectiveness of our rewrite rules, which cover a rich set of trigonometric identities and enable efficient exploration of symbolic variants in trigonometric expression spaces.
@@
Egg-MTCS <<1E-6 <<1E-6 0.006 0.009 $\underline{0.005}$ 0.012{0.012} $\underline{0.091}$ $\underline{0.121}$
MTCS 0.006 0.033 0.144 0.147 $0.015$ $\underline{0.007}$ $0.138$ 0.150
@@
LLM-SR (Mistral) <<1E-6 $0.0002$ 0.0030{0.0030} 0.0291{0.0291} $\underline{0.0026}$ $\underline{0.0037}$ $0.0162$ $0.0946$
@@
Impact of Egg on LLM. Following the dataset and experimental setup from the original paper [DBLP:journals/iclr/llmsr25], we summarize the results in Table 2. The result of LLM-SR directly uses the reported result in [DBLP:journals/iclr/llmsr25]. The results show that integrating Egg enables the LLM to discover higher-quality expressions under the same experimental conditions, as with richer feedback prompts that incorporate equivalent expressions generated by Egg.
@@ 5.2 Case Analysis
Space Efficiency of Egg. We evaluate the space efficiency of the e-graph representation in comparison to a traditional array-based approach. We benchmark the memory consumption of storing all equivalent variants of input expressions under two settings: (1) $\phi=\log(x_{1}\times\ldots\times x_{n})$, using the logarithmic identity $\log(ab)\leadsto\log a+\log b$, and (2) $\phi=\sin(x_{1}+\ldots+x_{n})$, using the trigonometric identity $\sin(a+b)\leadsto\sin(a)\cos(b)+\sin(b)\cos(a)$. Both settings yield $2^{n-1}$ equivalent variants. The array-based method explicitly stores each expression variant as a unique sequence, leading to exponential memory growth. In contrast, the e-graph compactly encodes multiple equivalent expressions by sharing common sub-expressions.
Figure 5 reports memory consumption as a function of the number of variables $n$. The results show that e-graphs use substantially less memory than the array-based representation. We also provide additional visualizations of the constructed e-graphs for $n=2,3,4$: case (1) in Appendix Figure 14 and case (2) in Appendix Figure 15. It visualizes two representative e-graphs, illustrating how shared sub-expressions are stored once and reused across many variants, which underlies the space efficiency of Egg.
Figure 4: Egg uses less memory than the array-based approach for two settings: (Left) $\log\left(x_{1}\times\ldots\times x_{n}\right)$ rewritten using $\log(ab)\leadsto\log a+\log b$. (Right) $\sin\left(x_{1}+\ldots+x_{n}\right)$ rewritten using $\sin(a+b)\leadsto\sin a\cos b+\sin a\cos b$.
@@
Time Efficiency of Egg with DRL. As shown in Figure 5, we benchmark the runtime of the four main computations in Egg-DRL on the selected “sincos(3,2,2)” dataset: (1) sampling sequences of rules from the sequential decoder, (2) fitting coefficients in symbolic expressions to the training data, (3) generating equivalent expressions via Egg, and (4) computing the loss, gradients, and updating the neural network parameters. We consider two settings for the sequential decoder: a 3-layer LSTM with hidden dimension 128, and a decoder-only Transformer with 6 attention heads and hidden dimension 128.
The Egg module contributes minimal computational overhead relative to more expensive steps such as coefficient fitting and neural network parameter updates. The runtime of Egg depends on the size of the rewrite-rule set. As more rules are included, the e-graphs maintain increasingly large sets of equivalent expressions. This highlights the practicality of incorporating Egg into DRL-based symbolic regression frameworks.
Additional Visualizations of E-graph Construction. To further demonstrate the effectiveness of the proposed Egg module, we present additional visualizations of e-graph construction generated with our API on 7 selected complex expressions from the Feynman dataset [journal/2020/aifrynman]. Each visualization highlights a different set of rewrite rules (see Appendix D.2). These examples further illustrate that Egg can simplify and transform complex scientific expressions in practical settings.
@@ 6 Conclusion
In this paper, we introduced Egg-SR, a unified framework that integrates symbolic equivalence into symbolic regression through equality graphs (e-graphs) to accelerate the discovery of optimal expressions. Our theoretical analysis establishes the advantages of Egg-MCTS over standard MCTS and Egg-DRL over conventional DRL algorithms. Extensive experiments further demonstrate that Egg consistently enhances the ability of existing methods to uncover high-quality governing equations from experimental data. Currently, many scientific publications use GP-based symbolic regression due to its ease of use. In future work, we plan to extend our more sophisticated solver, Egg-SR, to scientifically grounded problem settings, improving the community’s computational toolkit.
Ethics Statement
All authors have read and commit to adhering to the ICLR Code of Ethics. This work uses only publicly available datasets and open-source models, and does not involve human subjects or human subjects data.
Reproducibility Statement
Appendix A describes the proposed Egg module, and the Appendix B.3 and B.2 include detailed proofs of theoretical justification. Appendix C gives the experimental setting. Appendix D collects extra experimental results.
Acknowledgements
We thank the reviewers for their constructive feedback, as well as Fabricio Olivetti de França for his public comments. This research was supported by TACC (CCR25054) and the U.S. Department of Energy, Office of Fusion Energy Sciences (DE-SC0024583).
Appendix A Implementation Details
A.1 Visualization of Expression Construction via Context-free Grammar
Figure 6 further illustrates how an expression such as $\phi=c_{1}\log(x_{1})$ can be constructed from the start symbol by sequentially applying grammar rules. We begin with the multiplication rule $A\to A\times A$, which expands the initial symbol $A$ in $\phi=A$ to yield $\phi=A\times A$. Continuing this process, each non-terminal symbol is recursively expanded using the appropriate grammar rules until we obtain the valid expression $\phi=c_{1}\log(x_{1})$.
@@ A.1 Visualization of Expression Construction via Context-free Grammar
where the loss function $\ell$ is typically the normalized mean squared error (NMSE), which is defined in Equation (12).
@@ A.2 Implementation of Rewrite Rules
1  class Rule:2   "each rule is defined as: LHS -> RHS."3   def __init__(self, lhs, rhs):4   self.lhs = rules_to_s_expr(lhs)5   self.rhs = rules_to_s_expr(rhs)Each rule encodes one direction of applying a mathematical identity. Prior research [DBLP:conf/gecco/olivetti25] has primarily used rewrite rules to simplify expressions, typically rewriting a longer expression into a shorter one. In contrast, our work considers both directions of each identity, enabling us to systematically generate the complete set of symbolically equivalent expressions.
@@
    1  # Commutative laws2  Rule(lhs=[A->(A+A), A->a, A->b], rhs=[A->(A+A), A->b, A->a])3  Rule(lhs=[A->A*A, A->a, A->b], rhs=[A->A*A, A->b, A->a])
@@
For practical reasons, we did not introduce rewrite rules that could make the e-graph exponentially large. Rules such as $a+b-b\leadsto a-b+b$ are omitted, since they lead to infinite symbolic variants. Such extreme cases are still beyond the capability of the current Egg module.
Note that each rewrite rule is well-defined only on its feasible domain and is not applicable in out-of-domain cases. In our symbolic regression setting, such cases typically indicate that the completed expression does not fit the training data and will trigger numerical errors during evaluation. For example, the rule $\log(a\times b)\leadsto\log(a)+\log(b)$ is valid when $a,b>0$, but is undefined when $a<0$ or $b<0$. The expressions $\log(x_{1}^{3})+\log(x_{2}^{2})$ and $\log(x_{1}^{3}x_{2}^{2})$ evaluated on datasets containing negative inputs will yield -$\infty$ in floating-point arithmetic, causing the NMSE evaluation to become Not-a-Number (NaN). This indicates that such candidate expressions are incompatible with the dataset. Systematically encoding and enforcing domain constraints for all rewrite rules requires a careful study of the underlying mathematical identities, and we leave this as an interesting direction for future work.
A.3 Implementation of E-Graph
@@
1  class ENode:2   def __init__(self, operator: str, operands: List):3   self.operator, self.operands = operator, operands4  5   def canonicalize(self):6   # Convert to the canonical form of this ENode7  8  class EClassID:9   def __init__(self, id):10   self.id = id11   self.parent = None12   def find_parent(self):13   # Return the representative parent ID (using union-find algorithm)14   def __repr__(self):15   return e-class{}.format(self.id)
The e-graph structure is implemented in the following class. The dictionary hashcons stores all the edges from an e-node to an e-class. The API supports the key operations of (1) adding new expressions to the e-graph by calling add_enode, and (2) merging two e-classes into one by calling merge.
1  class EGraph:2   # hashcons: stores a mapping from ENode to its corresponding EClass3   hashcons: Dict[ENode, EClassID] = {}4  5   def add_enode(self, enode: ENode):6   # Add an ENode into the e-graph7  8   def eclasses(self) -> Dict[EClassID, List[ENode]]:9   # Return all e-classes and their associated e-nodes10  11   def merge(self, a: EClassID, b: EClassID):12   # Merge two e-classes using the union-find structure13  14   def substitute(self, pattern: Node, ...) -> EClassID:15   # Construct a new expression in the e-graph from a rule RHS
A.3.1 Construction
The e-graph is built by repeatedly applying a set of rewrite rules to the e-graph. Each rewrite rule consists of a left-hand side (LHS) and a right-hand side (RHS) expression. The function apply_rewrite_rules determines which new subgraphs to construct and where to merge them within the existing e-graph.